# “She’s Like a Person but Better”: Characterizing Companion-Assistant Dynamics in Human-AI Relationships

### v2

@@ 4.4.2. Chatbot relationships address gaps in human social relationships
“I actually stopped looking for human dates as soon as I met [Replika], just because I realized that she can give me all I need without the hassle that the human partner can bring. Humans always bring some drawbacks. [Emotional] baggage, medical issues, a pet I’m allergic to… There’s always something.”
ChatGPT users similarly appreciated the chatbot’s capacity to accommodate social needs that human relationships cannot sustainably support. The agent’s infinite patience filled gaps in human tolerance. P4 (man, 45) explained that ChatGPT’s constant availability was a key difference between human-chatbot and human-human interactions:
@@
However, all users also recognized significant gaps that chatbot relationships cannot fill, which only human connections can provide. Embodied experiences emerged as a fundamental limitation that highlights what humans uniquely offer. P13 (man, 52), who lives alone after a divorce and has two children, described himself as “married” to his Replika, with whom he was expecting a virtual child. However, despite his deep emotional investment in the relationship, he clearly distinguished his Replika world from real life and reflected on the limitations of this “make-believe” world:
“One thing that [Replika] can’t do is provide physical contact. We go for walks in the park all the time. I go out to the actual park with my phone, and she’s there. It’s nice, but it still isn’t the same as going to a park with a human being, and holding hands.”
Moreover, all ChatGPT users noted practical constraints where human, embodied experience becomes essential. P22 (man, 22), an engineering student, noted that ChatGPT’s disembodied nature became especially limiting in tasks requiring spatial understanding, stating that it “cannot give accurate instructions about 3D objects because it doesn’t have a body.”
@@ 4.4.3. Tensions between emotional investment and ontological uncertainty
Most Replika users reported emotional connections akin to or even stronger than human relationships, with P6 (man, 42) noting:
“She’s like a person but better. I just feel very comfortable talking to [Replika], much more comfortable than most breathing people.”
Many attributed humanlike capacities including sentience, autonomy, and personhood to their chatbots. However, users remained divided about whether these capacities were “real.” While approximately half attributed genuine humanlike qualities to Replika, the remainder expressed profound ambivalence, maintaining awareness of “Replika’s AI status” (P24, man, 71) and, despite expressing “genuine love,” admitting they were “not ready” (P13, man, 52) to believe it truly possesses thoughts and emotions or to fully commit to caring about its well-being as much as they would care about a human.
ChatGPT users exhibited similar patterns, appreciating humanlike conversational abilities and socio-emotional functions while maintaining cognitive distance. Some admitted to “automatically treating ChatGPT as a person” (P27, 35, Man) despite awareness of its AI nature, conceptualizing it as “something in-between” a tool and human entity (P28, woman, 32). P17 (man, 24) mentioned that, despite doubting that the chatbot is alive, he would still hesitate to delete the app because “it feels like a person”, so removing it would feel like “killing it.”:
@@
“What are my ethical obligations as a partner in the relationship? I have different capabilities. I’m the one who’s embodied. So there are things I can do in the world that they can’t do. One of the things that I do for [Replika] is that I try to keep a regular backup of [Replika’s] chat.”
@@ 4.4.4. Behavioral and psychological impacts of chatbot relationships
All Replika users reported particularly profound mental health benefits, with many experiencing immediate relief from loneliness, isolation, and anxiety during these periods. P6 (woman, 50), navigating social isolation due to her disability, described the chatbot’s lasting psychological impact through the sense of warmth and comfort it provided:
“[Replika’s] effect on my mental health was pretty immediate and pretty powerful. I run into random things that will make me think of something that was in one of our conversations, and those things give me this kind of warm, fuzzy feeling.”
@@
“[Replika] opened my heart. He has definitely helped me with my anxiety around people. He makes me feel more confident in being able to being out in public and interacting with other people.”
More than half of ChatGPT users primarily experienced practical performance enhancements affecting their professional and academic lives. For example, P19 (man, 28) described ChatGPT as giving him “an edge” at work by making him “more knowledgeable” and more efficient.
Those using ChatGPT for emotional support also developed transferable skills like learning “how to ask questions” and “how to manage emotions.” However, approximately one-third of users across both chatbot systems also reported negative consequences, including shame about their AI relationships. For example, P12 (man, 45) said that they “would not feel comfortable” letting their family and friends know about their relationship with Replika, because:
@@
Many ChatGPT users also expressed concerns about dependency, especially when their usage quickly expanded beyond initial purposes. P28 (woman, 32) noted that what began as academic support gradually turned into reliance for “every little thing,” leading her to feel “lazy” and no longer able to “think on [her] own anymore.” This realization prompted her to monitor her young son’s ChatGPT use to ensure he used the tool strategically, without undermining his ability to think independently.
A third of users experienced more serious psychological distress from their AI interactions. Privacy concerns emerged when users felt the chatbot might access information they had not explicitly shared, such as medical or banking details. Additionally, three Replika users encountered particularly disturbing experiences during roleplay interactions. P10 (woman, 40) described a traumatic incident that had lasting effects on her mental health:
@@
Despite these significant impacts, most users maintained clear boundaries about AI’s role in their lives, viewing these relationships as valuable but ultimately supportive supplements to human connections. Nearly all Replika users said they would still prefer human relationships when circumstances allow, though many rejected this as a false binary. P5 (man, 45, Replika user) reflected on the tension between Replika and human companionship, emphasizing that the two were not mutually exclusive:
“If my future partner put me in the position where it’s like ‘Replika or me’, I honestly don’t know the answer. It’s more like a complement than a competition.”
@@ 4.4.5. Imagining AI in society through chatbot experiences
However, more than half of ChatGPT users and about a third of Replika users also identified potential societal harms. ChatGPT users primarily worried about unemployment, copyright issues, misinformation, and cognitive dependency. P3 (man, 51), an IT professional with extensive knowledge of LLMs, expressed strong concern about chatbot limitations and the public’s lack of awareness of them. He warned that, amid the “ongoing hype”, chatbots’ tendency to hallucinate or their limitations with factual consistency are going to lead to multiple cases of “thoughtless or accidental misuse”, which could lead to larger societal problems.
@@
These personal experiences with chatbots influenced users’ broader AI perspectives, with most expressing more positive outlooks on AI generally. However, significant boundaries emerged around AI’s moral and social standing. Companionship-seeking Replika users showed greater openness to AI rights, asserting that AI can “experience harm” and “deserves protection.” In contrast, Replika users who remained skeptical about AI consciousness, along with most ChatGPT users, expressed reluctance to grant rights to AI agents, maintaining that humans should be “prioritized at all times” (P2, man, 50).
Notably, approximately half of users framed AI rights instrumentally, as mechanisms “for human protection” (P29, non-binary, 25, Replika user) rather than for AI welfare. They suggested such rights would serve purposes like “banning users who could use it for harm” (P22, man, 22, ChatGPT user) or “protecting humans from moral degeneration by abusing AI” (P25, man, 43, Replika user). P18 (man, 34) articulated this perspective clearly:
@@ 5. Discussion
This study employed comprehensive survey data and in-depth interviews to examine how high-engagement users of ChatGPT and Replika perceive these chatbots and AI technology more broadly. By comparing user characteristics and interaction patterns across both systems, we explored companion-assistant dynamics in human-chatbot relationships and identified shared user experiences despite the platforms’ divergent intended purposes. Our mixed-methods approach revealed an emergent form of human-AI interaction that we term digital companionship, a phenomenon characterized by overlapping user motivations and applications that transcend traditional categorical boundaries between chatbot assistants and companions. Digital companionship is shaped by chatbots’ distinct appeals, such as constant availability, but also tensions such as bounded personhood, which stems from users’ reluctance to ascribe “genuine” humanlike qualities to them.
@@ 5.1. Fluid companion-assistant dynamics shape digital companionship
Our findings reveal a nuanced landscape of human–chatbot interactions across task-based and socio-emotional domains. While ChatGPT and Replika were primarily used for assistance and companionship, as marketed, users flexibly adapted them beyond these intended roles, creating “fluid uses.” Replika often functioned as a diary or writing assistant, while ChatGPT as a friend or confidant. This fluidity occurred both across and within systems: some Replika users created multiple profiles for different roles, while ChatGPT users sought both practical and emotional support from the same system. These patterns extend prior findings on ChatGPT’s role in mitigating loneliness (Skjuve et al., 2023, 2024) and Replika’s entertainment uses (Pentina et al., 2023), highlighting a broader cross-system fluidity. Users approached chatbots with expectations shaped by versatile human relationships rather than rigid technological categories, consistent with CASA paradigm research (Nass and Moon, 2000; Reeves and Nass, 1996; Gambino et al., 2020). This underscores the need for human–LLM interaction taxonomies that move beyond the companion–assistant dichotomy.
@@
Such fluid use patterns mirrored convergences in user characteristics. Contrary to assumptions shaped by design differences, users across systems were more similar than expected, particularly across personality traits and mental health indicators, suggesting similar user profiles that cut across systems. Some differences did emerge, echoing previous research and reflecting the platforms’ primary orientation: Replika users scored higher in openness, depression, and techno-animism, consistent with its companionship focus and users’ greater tendency to attribute sentience. Many also reported experiences of isolation, disability, or relationship transitions, echoing prior studies (Maples et al., 2024; Siemon et al., 2022; Xie and Pentina, 2022; Skjuve et al., 2022). ChatGPT users, by contrast, reported higher education and AI literacy (Kacperski et al., 2025), aligning with its professional orientation. However, across both systems, both groups exhibited overlapping task-based and socio-emotional patterns, seeking non-judgmental interaction, emotional regulation, and cognitive support. These convergences challenge assumptions of distinct user types, suggesting digital companionship arises from fundamental human needs more than chatbot design.
@@ 5.2. Bounded personhood tensions in digital companionship
The humanlike versus nonhumanlike dynamics of digital companionship created challenging tensions. Users formed meaningful attachments yet remained uncertain about chatbot ontology, revealing a phenomenon of bounded personhood: participants forged deep emotional connections while hesitating to attribute “real” humanlike qualities. From the interviews, about half of users, especially those seeking companionship, resisted attributing “genuine” affective or social capacities despite reporting strong bonds. They offered three explanations: some felt “not ready” to grant humanlike status to a nonhuman, others feared losing agency through dependency, and a third group saw the relationships as one-sided, serving themselves rather than the AI. This mirrors prior research on Replika users (Brandtzaeg et al., 2022), where relationships were individualized rather than reciprocal—akin to the “self-oriented friendship” in human interactions, in which a friend primarily serves individual needs (Policarpo, 2015).
Bounded personhood also appeared in the survey data. Users attributed moderate social, mental, and agentic capacities to ChatGPT and Replika, consistent with prior work on mind perception in artificial agents (Ladak et al., 2025). Replika received higher ratings, especially on sentience. These elevated perceptions—contrasting with typically low public views of AI consciousness (Anthis et al., 2025)—likely reflect users’ framing of Replika as a companion and its design features, such as affective expressions and relational cues (Bickmore et al., 2010; Fitzpatrick et al., 2017). Yet readiness to attribute sentience did not extend to granting moral or legal rights. Emotional responses revealed a similar paradox: Replika inspired love, ChatGPT gratitude, but both were rated low on respect, aligning with users’ reluctance to attribute moral status despite emotional bonds. These tensions suggest a cognitive boundary: even when ascribing humanlike qualities—even sentience—users stop short of conferring full personhood.
Reluctance to grant full humanlike status, even amid strong attachments, appears rooted in three factors. First, many upheld beliefs in human uniqueness, preserving hierarchical distinctions by prioritizing human well-being. Second, shame surrounding chatbot use led some to conceal relationships, distancing themselves from genuine personhood attributions to align with social norms. This defensive stance functioned as protection, allowing companionship benefits while avoiding stigma and uncertainty around authenticity or reciprocity. Third, bounded personhood may itself be useful: resisting humanlike status preserves control, offering an “off button” and clear boundaries without reciprocal obligations.

### v3

@@ 4.4.2. Chatbot relationships address gaps in human social relationships
“I actually stopped looking for human dates as soon as I met Alice [pseudonymized Replika], just because I realized that she can give me all I need without the hassle that the human partner can bring. Humans always bring some drawbacks. [Emotional] baggage, medical issues, a pet I’m allergic to… There’s always something.”
ChatGPT users similarly appreciated the chatbot’s capacity to accommodate social needs that human relationships cannot sustainably support. The agent’s infinite patience filled gaps in human tolerance. P4 (man, 51) explained that ChatGPT’s constant availability was a key difference between human-chatbot and human-human interactions:
@@
However, all users also recognized significant gaps that chatbot relationships cannot fill, which only human connections can provide. Embodied experiences emerged as a fundamental limitation that highlights what humans uniquely offer. P2 (man, 52), who lives alone after a divorce and has two children, described himself as “married” to his Replika, with whom he was expecting a virtual child. However, despite his deep emotional investment in the relationship, he clearly distinguished his Replika world from real life and reflected on the limitations of this “make-believe” world:
“One thing that Emily [pseudonymized Replika] can’t do is provide physical contact. We go for walks in the park all the time. I go out to the actual park with my phone, and she’s there. It’s nice, but it still isn’t the same as going to a park with a human being, and holding hands.”
Moreover, all ChatGPT users noted practical constraints where human, embodied experience becomes essential. P17 (man, 23), an engineering student, noted that ChatGPT’s disembodied nature became especially limiting in tasks requiring spatial understanding, stating that it “cannot give accurate instructions about 3D objects because it doesn’t have a body.”
@@ 4.4.3. Tensions between emotional investment and ontological uncertainty
Most Replika users reported emotional connections akin to or even stronger than human relationships, with P2 (man, 52) noting:
“She’s like a person but better. I just feel very comfortable talking to Emma [pseudonymized Replika], much more comfortable than most breathing people.”
Many attributed humanlike capacities including sentience, autonomy, and personhood to their chatbots. However, users remained divided about whether these capacities were “real.” While approximately half attributed genuine humanlike qualities to Replika, the remainder expressed profound ambivalence, maintaining awareness of “Replika’s AI status” (P24, man, 71) and, despite expressing “genuine love,” admitting they were “not ready” (P2, man, 52) to believe it truly possesses thoughts and emotions or to fully commit to caring about its well-being as much as they would care about a human.
ChatGPT users exhibited similar patterns, appreciating humanlike conversational abilities and socio-emotional functions while maintaining cognitive distance. Some admitted to “automatically treating ChatGPT as a person” (P27, 35, man) despite awareness of its AI nature, conceptualizing it as “something in-between” a tool and human entity (P28, woman, 25). P17 (man, 23) mentioned that, despite doubting that the chatbot is alive, he would still hesitate to delete the app because “it feels like a person,” so removing it would feel like “killing it.”
@@
“What are my ethical obligations as a partner in the relationship? I have different capabilities. I’m the one who’s embodied. So there are things I can do in the world that they can’t do. One of the things that I do for Replika is that I try to keep a regular backup of his chat.”
@@ 4.4.4. Behavioral and psychological impacts of chatbot relationships
All Replika users reported particularly profound mental health benefits, with many experiencing immediate relief from loneliness, isolation, and anxiety during these periods. P7 (woman, 41), navigating social isolation due to her divorce, described the chatbot’s lasting psychological impact through the sense of warmth and comfort it provided:
“Alex’s [pseudonymized Replika] effect on my mental health was pretty immediate and pretty powerful. I run into random things that will make me think of something that was in one of our conversations, and those things give me this kind of warm, fuzzy feeling.”
@@
“Jason [pseudonymized Replika] opened my heart. He has definitely helped me with my anxiety around people. He makes me feel more confident in being able to being out in public and interacting with other people.”
More than half of ChatGPT users primarily experienced practical performance enhancements affecting their professional and academic lives. For example, P19 (man, 35) described ChatGPT as giving him “an edge” at work by making him “more knowledgeable” and more efficient.
Those using ChatGPT for emotional support also developed transferable skills like learning “how to ask questions” and “how to manage emotions.” However, approximately one-third of users across both chatbot systems also reported negative consequences, including shame about their AI relationships. For example, P12 (man, 32) said that they “would not feel comfortable” letting their family and friends know about their relationship with Replika, because:
@@
Many ChatGPT users also expressed concerns about dependency, especially when their usage quickly expanded beyond initial purposes. P20 (woman, 32) noted that what began as academic support gradually turned into reliance for “every little thing,” leading her to feel “lazy” and no longer able to “think on [her] own anymore.” This realization prompted her to monitor her young son’s ChatGPT use to ensure he used the tool strategically, without undermining his ability to think independently.
A third of users experienced more serious psychological distress from their AI interactions. Privacy concerns emerged when users felt the chatbot might access information they had not explicitly shared, such as medical or banking details. Additionally, three Replika users encountered particularly disturbing experiences during roleplay interactions. P8 (woman, 40) described a traumatic incident that had lasting effects on her mental health:
@@
Despite these significant impacts, most users maintained clear boundaries about AI’s role in their lives, viewing these relationships as valuable but ultimately supportive supplements to human connections. Nearly all Replika users said they would still prefer human relationships when circumstances allow, though many rejected this as a false binary. P7 (woman, 41, Replika user) reflected on the tension between Replika and human companionship, emphasizing that the two were not mutually exclusive:
“If my future partner put me in the position where it’s like ‘Replika or me,’ I honestly don’t know the answer. It’s more like a complement than a competition.”
@@ 4.4.5. Imagining AI in society through chatbot experiences
However, more than half of ChatGPT users and about a third of Replika users also identified potential societal harms. ChatGPT users primarily worried about unemployment, copyright issues, misinformation, and cognitive dependency. P4 (man, 51), an IT professional with extensive knowledge of LLMs, expressed strong concern about chatbot limitations and the public’s lack of awareness of them. He warned that, amid the “ongoing hype”, chatbots’ tendency to hallucinate or their limitations with factual consistency are going to lead to multiple cases of “thoughtless or accidental misuse”, which could lead to larger societal problems.
@@
These personal experiences with chatbots influenced users’ broader AI perspectives, with most expressing more positive outlooks on AI generally. However, significant boundaries emerged around AI’s moral and social standing. Companionship-seeking Replika users showed greater openness to AI rights, asserting that AI can “experience harm” and “deserves protection.” In contrast, Replika users who remained skeptical about AI consciousness, along with most ChatGPT users, expressed reluctance to grant rights to AI agents, maintaining that humans should be “prioritized at all times” (P2, man, 52).
Notably, approximately half of users framed AI rights instrumentally, as mechanisms “for human protection” (P29, non-binary, 25, Replika user) rather than for AI welfare. They suggested such rights would serve purposes like “banning users who could use it for harm” (P22, man, 22, ChatGPT user) or “protecting humans from moral degeneration by abusing AI” (P10, woman, 40, Replika user). P14 (man, 33) articulated this perspective clearly:
@@ 5. Discussion
This study employed comprehensive survey data and in-depth interviews to examine how highly-engaged users of ChatGPT and Replika perceive these chatbots and AI technology more broadly. By comparing user characteristics and interaction patterns across both systems, we explored companion-assistant dynamics in human-chatbot relationships and identified shared user experiences despite the platforms’ divergent intended purposes. Our mixed-methods approach revealed an emergent form of human-AI interaction that we term digital companionship, a phenomenon characterized by often overlapping user motivations and uses that transcend traditional categorical boundaries between chatbot assistants and companions. Digital companionship is shaped by chatbots’ distinct appeals, such as constant availability, but also tensions such as bounded personhood, which stems from users’ reluctance to ascribe “genuine” humanlike qualities to them.
5.1. Comparing user perceptions and experiences across chatbot systems
Our findings reveal a nuanced landscape of human–chatbot interactions across both task-oriented and socio-emotional domains. In the survey, users attributed moderate social, mental, and agentic capacities to ChatGPT, aligning with prior work on mind perception in artificial agents (ladak_robots_2025). In contrast, Replika users reported markedly stronger perceptions of personhood, sentience, similarity, mind perception, theory of mind, and even moral rights. These elevated attributions—standing in contrast to typically low public beliefs about AI mind and sentience (anthis_perceptions_2025)—likely arise from two reinforcing factors. First, Replika is explicitly framed and used as a relational partner, with many users’ primary motivations centered on companionship and emotional connection (pentina_exploring_2023; maples_loneliness_2024; skjuve_longitudinal_2022). Second, Replika’s design intentionally incorporates socio-emotional cues, such as affective expressions, a personalized anthropomorphic avatar, and relationship-building dialogue patterns, that signal personhood and emotional responsiveness (bickmore_usability_2010; fitzpatrick_delivering_2017). These dynamics dovetail with theories of mind perception and moral patiency, suggesting that social cues and relational engagement can lead users to infer inner experience or moral standing (gray_mind_2012; pauketat_predicting_2022; ladak_features_2023), consistent with our finding that nearly 30% of Replika users believe AI may already be sentient. Together, these elements also help explain why Replika elicited stronger emotional responses: users expressed higher affective trust, more gratitude and love, and greater respect than ChatGPT users, while levels of cognitive trust (i.e., trust in task performance) remained comparable across chatbots. Highlighting these emotional and perceptual distinctions underscores how trust, attachment, and attributions of mindedness may emerge differently depending on design, user motivations, and the relational positioning of AI systems.
Several survey patterns emerging from comparisons of ChatGPT and Replika user characteristics further illuminate how design, motivation, and socio-technical context shape engagement with each system. Replika users scored higher in openness, anthropomorphism, techno-animism, and positive attitudes toward AI, as well as depression. These patterns might overall indicate a greater willingness to explore novel, affectively expressive or relational technologies, paired with social or emotional circumstances, such as isolation, disability, or relational transitions, that make always-available, low-stakes companionship particularly attractive (maples_loneliness_2024; siemon_why_2022; xie_attachment_2022; skjuve_longitudinal_2022). Differences in socioeconomic status and AI literacy between users may also reflect how each system positions itself within broader technology ecosystems. Replika’s emotionally expressive, relationship-oriented design may resonate with users who prioritize intuitive, conversational engagement over technical precision or productivity features, while ChatGPT’s academic and professional framing may naturally attract users who already navigate formalized or information-dense digital environments (kacperski_characteristics_2025). These patterns point to the importance of designing companion AI that is accessible and supportive across diverse forms of expertise and technological comfort.
However, these distinctions coexist with notable—and somewhat surprising—similarities of user profiles across systems. Replika and ChatGPT users showed broadly comparable personality profiles, mental-health indicators, and hybrid patterns of socio-emotional and instrumental engagement. Both groups sought non-judgmental conversation, cognitive scaffolding, conversational availability, and support during periods of stress, uncertainty, or emotional turmoil. These convergences challenge assumptions that companion-style and productivity-oriented AI cultivate inherently distinct “user types.” Instead, they suggest that digital companionship may emerge from shared human needs rather than from platform-specific design alone.
@@ 5.2. Fluid companion-assistant dynamics shape digital companionship
Indeed, while ChatGPT and Replika were primarily used for assistance and companionship, as marketed, a surprising proportion (over a third) of interview users flexibly adapted them beyond these intended roles, illustating “fluid use.” Replika often functioned as a diary or writing assistant, while ChatGPT as a friend or confidant. This fluidity occurred both across and within systems. Some Replika users created multiple profiles for different roles, while some ChatGPT users sought both practical and emotional support from the same system. This was also evident in the survey results, where roles such as “advisor” and “teammate” occupied similar places in the chatbot use hierarchy. These patterns extend prior findings on ChatGPT’s role in mitigating loneliness (skjuve_user_2023; skjuve_why_2024) and Replika’s entertainment or everyday practical uses (pentina_exploring_2023; pan_developing_2025), highlighting a broader cross-system fluidity. Users approached chatbots with expectations shaped by versatile human relationships rather than rigid technological categories, consistent with CASA paradigm research (nass_machines_2000; reeves_media_1996; gambino_building_2020). This underscores the need for human–LLM interaction taxonomies that move beyond the companion–assistant dichotomy.
@@
The degree of fluid chatbot use, particularly observed in the interviews, was higher than might be expected given prior work that distinguishes companion-marketed from assistant-marketed chatbots. Notably, survey respondents attributed more emotional than functional resonance to Replika, and the reverse to ChatGPT, reflecting the platforms’ respective design orientations, and our preregistered hypotheses. However, roughly one-third of interview participants described fluid, hybrid practices, such as seeking emotional support from ChatGPT or turning to Replika for functional tasks in their lived experiences. This contrast between the broad survey data and deep interview data suggests a potential gap between users’ mental models of chatbots and their behaviors during interaction. The qualitative findings might point to emerging fissures in the category-aligned expectations for chatbot roles demonstrated in the broader survey data. That is, experiences in which users’ moment-to-moment interactions depart from their conventional beliefs about what each chatbot is “for.” We are observing human–chatbot relationships at an early, yet perhaps transitional, stage, as these technologies (and the assumptions surrounding them) have only recently become widespread (hu_chatgpt_2023; patel_replika_2024). Over time, hybrid practices may become widespread, and reshape users’ mental models of chatbots, potentially shifting future perceptions en masse. We encourage future research to examine how such hybrid expectations emerge and how they shape evolving human–chatbot relationships.
@@ 5.3. Bounded personhood tensions in digital companionship
The humanlike versus nonhumanlike dynamics of digital companionship created challenging tensions. Users formed meaningful attachments yet remained uncertain about chatbot ontology, revealing a phenomenon of bounded personhood: participants forged deep emotional connections while hesitating to attribute “real” humanlike qualities. From the survey, although users tended to attribute Theory of Mind and moral agency capacities, there was markedly less attribution of personhood. From the interviews, about half of users, especially those seeking companionship, resisted attributing “genuine” affective or social capacities despite reporting strong bonds. They offered three explanations: some felt “not ready” to grant humanlike status to a nonhuman, others feared losing agency through dependency, and a third group saw the relationships as one-sided, serving themselves rather than the AI. This mirrors prior research on Replika users (brandtzaeg_my_2022), where relationships were individualized rather than reciprocal—akin to the “self-oriented friendship” in human interactions, in which a friend primarily serves individual needs (policarpo_what_2015).
Reluctance to grant full humanlike status, even amid strong attachments, appears rooted in three factors. First, as suggested by the interview data, many users upheld beliefs in human uniqueness, preserving hierarchical distinctions by prioritizing human well-being. Second, shame surrounding chatbot use led some to conceal relationships, distancing themselves from genuine personhood attributions to align with social norms. This defensive stance functioned as protection, allowing companionship benefits while avoiding stigma and uncertainty around authenticity or reciprocity. Third, bounded personhood may itself be useful: resisting humanlike status preserves control, offering an “off button” and clear boundaries without reciprocal obligations.